OpenStudy (anonymous):

write as single trigonometric function: sin(90degrees-theta)/cot(90degrees-theta) PLEASE I KNOW THE ANSWER I JUST NEED SOMEONE TO EXPLAIN IT

OpenStudy (misty1212):

\[\sin(90-x)=\frac{a}{h}=\cos(x)\]

OpenStudy (misty1212):

same with \[\cot(90-x)=\frac{b}{a}=\tan(x)\]

OpenStudy (nikato):

You should know by cofunction identities That sin(90-theta)= cos theta And cot(90-theta)=tan theta http://www.mathwords.com/t/trig_identities.htm So you can rewrite as cos(theta) ---------- tan(theta)

OpenStudy (anonymous):

but the her said the answer is just cot theta

OpenStudy (anonymous):

teacher

OpenStudy (misty1212):

the teacher is mistaken

OpenStudy (misty1212):

unless there is a typo in the question

OpenStudy (anonymous):

oh yes my apologies its cos not cot

OpenStudy (anonymous):

i understand now